OpenStudy (anonymous):

64x^2-80x+25 factor as perfect square trinomial whereever possible

OpenStudy (compassionate):

(8x - 5^2)

OpenStudy (compassionate):

Do you understand why?

OpenStudy (mathstudent55):

a^2 - 2ab + b^2 = (a - b)^2 Here, a^2 = 64x^2, so a = 8x Also, here b^2 = 25, so b = 5

OpenStudy (anonymous):

I do not think that one is correct

OpenStudy (compassionate):

(8x - 5)^2 is correct